### v1

# BarrierBench : Evaluating Large Language Models for Safety Verification in Dynamical Systems

Abstract
Safety verification of dynamical systems via barrier certificates is central to ensuring correctness in autonomous and safety-critical applications. Synthesizing such certificates requires discovering mathematical functions that characterize inductive state invariants which provably separate safe and unsafe regions. However, existing approaches struggle with computational scalability, depend on careful template design, and often rely on exhaustive or incremental searches over function spaces. Moreover, these approaches demand significant manual ingenuity and mathematical sophistication in constructing the search infrastructure: selecting template structures, choosing solvers, tuning hyperparameters for data-driven methods, and designing effective sampling procedures. A successful synthesis of barrier certificates therefore requires both a deep understanding of system and control theory and practical experience with these techniques. This expertise has traditionally been passed among practitioners through natural language rather than being formalized mathematically.
This observation raises a natural question: can the linguistic and analogical reasoning that experts use informally be captured and operationalized by language models? Motivated by this idea, we present a Large Language Model (LLM) agentic framework for barrier certificate synthesis that leverages natural language reasoning to propose, refine, and validate candidate certificates. The framework integrates LLM-driven template discovery with SMT-based verification and supports barrier-controller co-synthesis for controlled systems, ensuring mathematical compatibility between safety certificates and feedback control laws.
To evaluate this capability, we introduce BarrierBench, a benchmark of 100 dynamical systems spanning linear, nonlinear, discrete-time, and continuous-time settings, including 68 controlled systems that require barrier–controller co-synthesis. Our experiments assess not only the effectiveness of LLM-guided barrier synthesis but also the utility of retrieval-augmented generation (RAG) and agentic coordination strategies in improving its reliability and performance. Across these tasks, the framework achieves more than 90% success in generating valid certificates and demonstrates structural diversity ranging from simple quadratics to high-order coupled polynomials. By releasing BarrierBench and the accompanying toolchain, we aim to establish a community testbed for advancing the integration of language-based reasoning with formal verification in dynamical systems. The benchmark is publicly available at https://hycodev.com/dataset/barrierbench.
footnotetext: 1 Isfahan University of Technology, Iran. (Email: {taheri.a, taban.a}@ec.iut.ac.ir) 2 Max Planck Institute for Software Systems, Germany. (Email: sadegh@mpi-sws.org) 3 University of Colorado Boulder, USA. (Email: ashutosh.trivedi@colorado.edu)
Keywords Barrier Certificates $\cdot$ LLMs $\cdot$ Safety Verification $\cdot$ Benchmarking and Reproducibility
1 Introduction
This century has witnessed an increasingly sophisticated integration of digital software with physical infrastructure, interconnected through wired and wireless networks. These cyber-physical systems (CPS) operate so seamlessly and intuitively that it is often easy to overlook the complexity beneath their surface: their success depends on intricate control logic whose correctness requires a deep understanding of both the underlying physics and the discrete switching logic that governs their operation. As a result, while technological practice continues to deliver critical infrastructures—such as autonomous vehicles, implantable medical devices, and smart grids—the task of control engineers to ensure their safety remains a Sisyphean endeavor amid ever-growing system complexity. In this paper, we address this enduring challenge by exploring how large language models can assist in the synthesis and verification of safety guarantees for complex CPS.
Safety Verification.
Safety verification of dynamical systems is a canonical challenge in the design of safety-critical CPS. Since safety verification is computationally undecidable even for CPS with simple piecewise-constant dynamics (asarin1998achilles), achieving provable safety through fully automatic procedures remains out of reach. Consequently, a range of techniques has been developed to construct safety proofs, including symbolic state-space exploration, statistical guarantees, and abstraction-based methods. A particularly effective and scalable alternative is the synthesis of barrier certificates (prajna2004safety; ames2019control; jagtap2020formal), which serve as functional analogs of inductive state invariants and characterize the separation between safe and unsafe regions of the state space. With origins in Lyapunov’s classical theory of stability (lyapunov1966stability; khalil2002nonlinear), barrier certificates extend its principle from proving convergence to demonstrating invariance. Their formulation as a search for a single scalar function subject to inequality constraints makes them amenable to optimization and satisfiability-based methods, contributing to their popularity in both theory and practice. We focus on the barrier-certificate-based approach to proof discovery as a promising foundation for integrating formal reasoning with language-model-guided synthesis.
Barrier Certificates.
Formally, a barrier certificate is a scalar function whose level sets define an inductive invariant that separates unsafe regions from all trajectories starting from admissible initial conditions. By providing formal guarantees over infinite time horizons, barrier certificates have become central to ensuring correctness in both linear and nonlinear dynamical systems. Traditional approaches to barrier synthesis are predominantly optimization-based. Sum-of-squares and semidefinite programming methods formulate the search as a convex optimization problem over polynomial templates (prajna2007framework; kordabad2025sum), yielding provable correctness but suffering from exponential complexity growth with system dimension and being restricted to low-degree polynomial classes. To improve generality, counterexample-guided inductive synthesis frameworks (see the work by (dawson2023safe) for a comprehensive review) extend these techniques using SMT solvers to iteratively refine candidate barriers (ravanbakhsh2015counter), whereas neural and data-driven variants employ deep architectures to approximate more expressive certificate functions (abate2024safe; neustroev2025neural; abate2021fossil). Although such methods broaden the space of representable barriers, they introduce new challenges—instability in training, reliance on finite samples, reduced interpretability, and a persistent need for careful template or architecture design. Consequently, successful synthesis continues to depend heavily on expert intuition: selecting suitable function classes, solvers, and hyperparameters. These choices often encode informal reasoning that remains difficult to capture or automate within existing algorithmic frameworks. This paper investigates whether large language models (LLMs) can capture and operationalize the linguistic and analogical reasoning that practitioners employ when constructing barrier certificates.
Language-Model-Guided Synthesis.
We present an LLM-guided synthesis framework for barrier certificates that leverages natural-language reasoning to propose, refine, and validate candidate certificates. The framework integrates LLM-driven template discovery with SMT-based formal verification and extends naturally to barrier–controller co-synthesis for controlled systems. By incorporating verification feedback into subsequent LLM prompts, the framework enables dynamic exploration of barrier certificate templates and coefficients, often discovering novel mathematical forms that traditional fixed-template approaches (prajna2007framework; kong2013exponential; ames2019control; abate2021fossil; kordabad2025sum) might overlook. LLMs guide the synthesis process through reasoning about system properties, safety requirements, and mathematical structures, systematically analyzing local safety properties before integrating them into global certificates. This integration provides both the expressiveness and adaptability needed for complex systems while preserving the mathematical rigor essential for safety-critical applications.
Figure 1: Architecture of the agentic framework integrating Retrieval, Synthesis, and Verifier agents for LLM-guided barrier certificate synthesis.
To operationalize this process, we design an agentic architecture that emulates the workflow of expert control designers. Human designers typically draw on prior experience with similar dynamical systems to select suitable barrier templates, solver strategies, and proof heuristics. Analogously, our framework instantiates three collaborating agents: a Barrier Retrieval Agent that searches a database of previously solved systems to identify analogous examples; a Barrier Synthesis Agent that proposes candidate barrier certificates inspired by retrieved analogs; and a Barrier Verifier Agent that evaluates these candidates in two stages—a lightweight, data-driven check followed by formal verification via an SMT solver. Feedback from the Verifier Agent is fed back to the Synthesis Agent, forming a closed-loop interaction among agents bounded by a global timeout. This design enables iterative refinement of candidate certificates and efficient use of external formal tools. Our study investigates the following research questions:
  1. RQ1:
    Are state-of-the-art large language models already capable of independently designing valid barrier certificates for safety verification tasks?
  2. RQ2:
    Does knowledge of barrier certificates for similar systems improve synthesis performance, and can retrieval-augmented generation (RAG) facilitate this transfer?
  3. RQ3:
    Can integrating specialized agents—retrieval, synthesis, and verification—into an agentic framework that interfaces with external solvers improve the efficiency and reliability of LLM-guided barrier synthesis?
BarrierBench.
To evaluate these questions, we introduce BarrierBench, a benchmark of 100 dynamical systems spanning linear, nonlinear, discrete-time, and continuous-time settings, including 68 controlled systems that require barrier–controller co-synthesis. Across these tasks, the LLM framework achieves more than 90% success in generating valid certificates and exhibits structural diversity ranging from simple quadratics to high-order coupled polynomials. By releasing BarrierBench and the accompanying toolchain, we aim to establish a community testbed for advancing the integration of language-based reasoning with formal verification in dynamical systems.
2 Barrier Certificates
A dynamical system is defined as a tuple $\mathcal{S}=(X,f,X_{I},X_{U})$, where $X\subseteq\mathbb{R}^{n}$ is the state space, $f:X\rightarrow\mathbb{R}^{n}$ is a vector field, $X_{I}\subseteq X$ is the set of initial states, and $X_{U}\subseteq X$ is the set of unsafe states with $X_{I}\cap X_{U}=\emptyset$. For continuous-time systems, the evolution is governed by
$\dot{x}(t)=f(x(t)),\quad x(0)\in X_{I},$ (1)
where $x(t)\in X$ denotes the state at time $t\geq 0$. For discrete-time systems, the evolution is given by
$x[k+1]=f(x[k]),\quad x[0]\in X_{I},$ (2)
where $k\in\mathbb{N}$ is the discrete time step.
Definition 1 (Safety Specification).
Given a dynamical system $\mathcal{S}=(X,f,X_{I},X_{U})$, $\mathcal{S}$ is safe if trajectories starting in $X_{I}$ remain in $X\setminus X_{U}$ for all time.
Definition 2 (Barrier Certificate).
A barrier certificate is a scalar function $B:X\rightarrow\mathbb{R}$ satisfying:
$B(x)$ $\leq 0,$ $\forall x\in X_{I},$ (3)
$B(x)$ $>0,$ $\forall x\in X_{U},$ (4)
$\nabla B(x)\cdot f(x)$ $<0,$ $\forall x\in X\text{ s.t. }B(x)=0\text{ (continuous-time)},$
$\text{ or }B(f(x))-B(x)$ $\leq 0,$ $\forall x\in X\text{ (discrete-time)}.$ (5)
Here, $\nabla B(x)\cdot f(x)$ denotes the Lie derivative $\mathcal{L}_{f}B(x)$ of a continuously differentiable function $B(x)$ along the vector field $f(x)$, defined as $\mathcal{L}_{f}B(x)=\sum_{i=1}^{n}\frac{\partial B}{\partial x_{i}}f_{i}(x)$.
Theorem 1 (Safety via Barrier Certificates (prajna2004safety)).
If there exists a barrier certificate $B(x)$ for $\mathcal{S}$, then the system $\mathcal{S}$ is safe.
The above definitions and theorem are stated for systems without control inputs. If the system has control input $u\in U$, the objective is to design a state feedback that gives the input $u$ as a function of state $x$, which makes the system safe. For this, the conditions of barrier will be quantified existentially over the input.
Input : System $(X,f,X_{I},X_{U})$; max iterations $K$; refinements $R$; input set $U$ for controlled systems
Output : Barrier $B^{*}$, controller $C^{*}$ for controlled systems, or failure
Initialize: best_score $\leftarrow 0$; $B^{*},C^{*}\leftarrow\text{null}$
for $k\leftarrow 1$ to $K$ do
 $\text{example}\leftarrow\textsc{RetrievalAgent}(f,X_{I},X_{U},\mathcal{D})$; $(B_{k},C_{k})\leftarrow\textsc{SynthesisAgent}(\text{example},\textsf{feedback}_{k})$
[…]
    continue
   end if
 $f^{\prime}\leftarrow f$ with $C_{k}$; $(\textsf{valid}_{k},\textsf{feedback}_{k})\leftarrow\textsc{VerifierAgent.FormalCheck}(B_{k},f^{\prime})$
 if $\textsf{valid}_{k}$ then
[…]
   end if
 for $r\leftarrow 1$ to $R$ do
    $(B_{k}^{r},C_{k}^{r})\leftarrow\textsc{SynthesisAgent.refine}(\textsf{feedback}_{k},r)$
[…]
       continue
      end if
    $f^{\prime}\leftarrow f$ with $C_{k}^{r}$; $(\textsf{valid}_{k}^{r},\textsf{feedback}_{k}^{r})\leftarrow\textsc{VerifierAgent.FormalCheck}(B_{k}^{r},f^{\prime})$
    if $\textsf{valid}_{k}^{r}$ then
[…]
      end if
   end for
end for
return Best partial solution
Algorithm 1 Agentic Barrier Certificate Synthesis
3 Agentic Framework for Barrier Synthesis
Our synthesis framework employs a multi-agent architecture in which specialized LLM-powered agents collaborate to discover barrier certificates. The framework comprises three core agents operating within an iterative pipeline: (1) a Barrier Retrieval Agent that identifies relevant historical solutions; (2) a Barrier Synthesis Agent that generates candidate barrier certificates; and (3) a Barrier Verifier Agent that verifies candidates and provides feedback for progressive refinement. This iterative process overcomes limitations of conventional approaches by enabling intelligent template selection and adaptive exploration of diverse mathematical structures through verification-guided synthesis.
Framework Pipeline.
The framework $\mathcal{F}$ takes as input a dynamical system $(X,f,X_{I},X_{U})$, with optional control input $U$ for controlled systems. A persistent database stores previously verified barrier certificates, which the Retrieval Agent accesses to accelerate synthesis on similar problems. The pipeline consists of four stages: (1) historical retrieval, (2) LLM-based synthesis, (3) two-stage verification, and (4) iterative refinement guided by feedback. The system performs up to $K$ main iterations, each exploring new template structures, and up to $R$ refinements per iteration to adjust coefficients or terms. The process terminates upon discovery of a valid barrier certificate or exhaustion of all attempts. If no valid certificate is found, the framework returns the candidate maximizing the proportion of satisfied conditions. Figure 1 illustrates the overall architecture of $\mathcal{F}$.
Barrier Retrieval Agent.
This agent analyzes the input system and retrieves relevant examples from the database through a two-stage process. First, it performs feature-based retrieval using automatically extracted system attributes such as dimension, time domain, linearity, and topology. Candidate problems matching all features are shortlisted. Second, an LLM-based reasoning stage ranks these candidates for structural similarity, providing the most compatible example to the Synthesis Agent while noting that adaptation may be required. When no match is found, the agent instructs the Synthesis Agent to proceed independently. As shown in Section 4.4, this retrieval step significantly accelerates convergence by enabling frequent first-try success.
Barrier Synthesis Agent.
This agent leverages the reasoning abilities of LLMs to analyze system dynamics and propose barrier certificates that satisfy the safety conditions (3)–(5). During the first main iteration ($k=1$), it incorporates relevant historical examples from the Retrieval Agent. For later iterations and refinements, it integrates feedback from the Verifier Agent to modify structure and coefficients. For systems requiring control synthesis, the agent jointly produces both the barrier certificate and controller expressions to ensure compatibility between safety and control objectives.
Barrier Verifier Agent.
The Barrier Verifier Agent conducts a two-stage evaluation of candidate barriers. In the first stage, it performs a sample-based validation across $N$ points drawn from $X_{I}$, $X_{U}$, and $X$, where $N$ is selected to balance computational efficiency with adequate state space coverage; we use $N=5000$ in our experiments to achieve this trade-off, discarding invalid barriers to avoid unnecessary SMT solver calls. In the second stage, it performs formal verification using SMT solvers such as Z3 (demoura2008z3), Yices (dutertre2006yices), and cvc5 (barbosa2022cvc5), checking the unsatisfiability of negated barrier conditions. The agent selects solvers adaptively based on problem structure, applies symbolic conversions or Taylor approximations when necessary, and handles verification failures through solver switching or adaptive timeout adjustments.
Iterative Refinement Mechanism.
When verification fails, the Verifier Agent provides the violated conditions and counterexamples as feedback to the Synthesis Agent, which uses them to refine the barrier. In early refinements ($r\leq 2$), the agent adjusts coefficients while preserving structure; in later refinements ($r>2$), it performs structural modifications such as adding or removing coupling terms while maintaining polynomial degree. This strategy balances exploitation (fine-tuning promising templates) and exploration (structural variation across iterations). The Verifier Agent logs all intermediate results, enabling cumulative feedback that helps the Synthesis Agent learn from previous attempts. As shown in Section 4.4, this mechanism substantially increases success rates, with a large number of cases solved within one or two main iterations.
4 Barrierbench: Experiments and Case Studies
The complete set of prompts used in our framework are included in the Appendix A.
4.1 Experimental Setup
The framework operates with carefully chosen hyperparameters validated across various systems. We set maximum main iterations $K=5$ and maximum refinements per iteration $R=4$. The LLM utilizes Claude Sonnet 4 and ChatGPT-4o with engineered prompts designed to synthesize mathematical expressions while maintaining formal correctness. All experiments use Z3 version 4.15.1, cvc5 version 1.3.1, and Yices version 2.6.4 as the SMT solvers for formal verification.
4.2 Case Studies
Table 1: Dataset characteristics of BarrierBench.
Characteristic Description
Dimensions 1D–8D
Initial set topology Ball, rectangle
Unsafe set topology Ball, rectangle, union of rectangles
System types Controlled, autonomous
Time domain Continuous-time, discrete-time
Dynamics types Linear, nonlinear
We evaluate our agentic barrier synthesis framework on BarrierBench, a benchmark comprising 100 dynamical systems. Our evaluation case studies encompasses continuous-time systems, discrete-time systems, and controlled systems requiring simultaneous barrier-controller co-synthesis, with a variety of complex nonlinear dynamics across dimensions from 1D to 8D. BarrierBench includes 68 controlled systems for which our framework demonstrates the simultaneous synthesis of barrier certificates and controllers parameters, which together establish safety guarantees for closed-loop dynamics. The characteristics of our benchmark are shown in Table 1.
4.3 Qualitative Results
In this section, we qualitatively review the results of some of the benchmarks synthesized by our framework, for more case studies you can check https://hycodev.com/dataset/barrierbench.
  • Continuous-Time Systems. The 4D system $\dot{x}_{1}=-x_{1}+4.5,\dot{x}_{2}=x_{1}-x_{2}-3,\dot{x}_{3}=-0.5x_{3},\dot{x}_{4}=-0.3x_{4}$ with state space $\mathbb{R}^{4}$ and polynomial degree of 1 has rectangular initial set $1.75,2.25]\times[1.75,2.25]\times[-0.1,0.1]\times[-0.1,0.1]$ and rectangular unsafe region $9,10]\times[9,10]\times[-5,5]\times[-5,5]$. The framework generates the quadratic barrier $B(x)=(x_{1}-4.5)^{2}+(x_{2}-1.5)^{2}+x_{3}^{2}+x_{4}^{2}-25$ of degree 2.
  • Discrete-Time Systems. The 2D system $x_{1}[k+1]=x_{1}[k]+0.01(-100x_{1}[k]-x_{2}[k]),x_{2}[k+1]=x_{2}[k]+0.01(x_{1}[k]-100x_{2}[k])$ with state space $\mathbb{R}^{2}$ and polynomial degree $d_{f}=1$ has rectangular initial set $0.1,0.4]\times[0.1,0.55]$ and rectangular unsafe region $0.45,0.5]\times[0.6,1.0]$. The framework generates the quadratic barrier $B(x)=x_{1}^{2}+x_{2}^{2}-0.5$ of degree 2.
  • Controlled Systems with Barrier-Controller Co-Synthesis. The 4D system $\dot{x}_{1}=x_{2}+u_{0},\dot{x}_{2}=-0.95\sin(x_{1})-0.02x_{2}+0.01x_{3}+u_{1},\dot{x}_{3}=-0.1x_{3}+0.02x_{1}+u_{2},\dot{x}_{4}=-0.03x_{4}+0.01x_{2}+u_{3}$ with state space $\mathbb{R}^{4}$, control space $\mathbb{R}^{4}$, and non-polynomial dynamics has initial ball with center $(0,0,0,0)$ and radius 0.3, and unsafe set as the complement of the ball with center $(0,0,0,0)$ and radius 2.5. The framework synthesizes controllers $u_{0}=-3x_{1}-1.5x_{2},u_{1}=-3x_{2}+0.8\sin(x_{1}),u_{2}=-3x_{3}-0.05x_{1},u_{3}=-3x_{4}-0.02x_{2}$ with quadratic barrier $B(x)=x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{4}^{2}-4.0$ of degree 2.
4.4 Ablation Study
In this section, we compare our full pipeline against a baseline that uses only a single LLM prompt without any framework support to evaluate our framework. Moreover, to evaluate the individual contributions of our framework components, we conduct ablation experiments. Table 2 presents the comparison between our framework and a single LLM prompt based on the number of solved problems in BarrierBench.
Table 2: Comparison of baseline (single-prompt) and agentic framework on BarrierBench.
Approach Claude Sonnet 4 ChatGPT-4o
Success Rate Solved Success Rate Solved
Baseline (Single Prompt) 41% 41/100 17% 17/100
Full Framework 90% 90/100 46% 46/100
Improvement +49% +49 +29% +29
The baseline approach, which relies on a single prompt LLM generation without any barrier retrieval from the historical database, iterative refinement, or failure-guided refinement, successfully solves only 41 problems (41%) with Claude Sonnet 4 and only 17 problems (17%) with ChatGPT-4o. In contrast, our full framework achieves substantial improvements on both models: 90% success rate with Claude Sonnet 4 and 46% success rate with ChatGPT-4o, representing improvements of +49 and +29 percentage points, respectively. This performance gap validates the significant role of our framework components and provides a negative answer to RQ1. The baseline’s limited success rates demonstrate that state-of-the-art LLMs cannot independently design valid barrier certificates without structured framework support. To further analyze the distribution of synthesized valid barriers across main iterations and refinements, Table 3 presents a detailed breakdown showing when and how problems are solved within our framework.
Table 3: Number of problems solved at each iteration–refinement combination. The index of II shows the number of iterations and the index of $R$ shows the number of refinements. For example, the number under $I_{1}$$R_{0}$ denotes the number of cases solved in first-try without refinement, leveraging prior database search for similar problems.
$\boldsymbol{R_{0}}$ $\boldsymbol{R_{1-2}}$ $\boldsymbol{R_{3-4}}$ Total
$\boldsymbol{I_{1}}$ 49 9 5 63 (70.0%)
$\boldsymbol{I_{2}}$ 10 3 2 15 (16.7%)
$\boldsymbol{I_{3+}}$ 4 2 6 12 (13.3%)
Total 63 (70.0%) 14 (15.6%) 13 (14.4%) 90 (100%)
Barrier Retrieval Impact.
As shown in Table 3, our barrier retrieval mechanism, which combines feature-based filtering and LLM selection accelerates convergence. Among the 90 successfully solved problems, 49 (54.4%) are solved in the first attempt without any refinements ($I_{1}-R_{0}$), compared to 41 total successes from the baseline. This indicates that our intelligent barrier retrieval not only increases the overall success rate but also accelerates solution discovery. Problems that require multiple iterations or fail entirely under the baseline approach are now solved by leveraging patterns from problems with matching properties, which provides a positive answer to RQ2.
Refinement Impact.
The refinement mechanism contributes substantially to the framework performance. As shown in Table 3, while 63 problems (70%) require no refinement ($R_{0}$ column), an additional 14 problems (15.6%) benefit from non-structural refinement ($R_{1-2}$), and 13 problems (14.4%) require structural adjustments ($R_{3-4}$). Within the first iteration alone, the refinement mechanism successfully synthesizes certificates for an additional 14 problems beyond the initial 49, bringing $I_{1}$ success to 63 solved problems. This demonstrates that even when the initial LLM-generated template requires adjustment, the refinement mechanism can improve coefficients and, when necessary, modify mathematical structures to satisfy all safety conditions.
Multi-Iteration Synthesis Impact.
As demonstrated in Table 3, for problems not solved in the first iteration, the framework’s ability to generate structurally different templates in later iterations proves essential, bringing additional 27 successful barrier synthesis. The second iteration ($I_{2}$) successfully solves 15 additional problems (16.7%), bringing the cumulative success rate to 86.7%. The remaining 12 problems (13.3%) require three or more iterations ($I_{3+}$), with half of these requiring structural refinement ($R_{3-4}$), indicating that more challenging problems often require exploration of alternative mathematical forms, as the proportion of problems requiring structural refinement increases in later iterations (from 7.9% in $I_{1}$ to 50% in $I_{3+}$).
The results of the ablation study provide a positive answer to RQ3, demonstrating the improvement in the efficiency and reliability of LLM-guided barrier synthesis through the agents interaction.
5 Related Work
Barrier Certificate Generation.
Traditional barrier synthesis has been dominated by optimization-based methods, particularly sum-of-squares (SOS) programming (prajna2007framework; ames2019control; wooding2025protect; kordabad2025sum). To overcome the scalability limitations of SOS, counterexample-guided inductive synthesis (CEGIS) frameworks such as FOSSIL (abate2021fossil) integrate neural function approximators with SMT-based verification, enabling synthesis of nonlinear certificates through iterative counterexample refinement. These approaches substantially expand the range of verifiable systems beyond fixed polynomial templates.
Application of barrier certificates to specifications beyond safety is also studied in the literature (lindemann2018control; jagtap2020formal; majumdar2024necessary). Data-driven variants extend barrier synthesis to systems with unknown or partially known dynamics (nejati2023formal). In the literature on data-driven barrier certificates, most existing methods are restricted to linear or control-affine system dynamics (see, e.g., (Jagtap2020CBCGP; Cohen2022; Lopez2022uCBF)). Moreover, these approaches often depend on known Lipschitz constants to ensure formal guarantees or are applicable only when the system dynamics are partially known. For example, Gaussian processes are used by (Wang2018CBF) and by (Jagtap2020CBCGP) to learn unknown components of nonlinear dynamics while assuming the control-affine part is known. Systems affected by unknown additive disturbances are studied by (Chekan2023UncertainConstraints). Similarly, (mathiesen2024data) assume that the deterministic component of the dynamics is accurately modeled, whereas the barrier-based safety control framework in of (mazouz2024data) presumes known noise distributions. Research addressing systems with fully unknown dynamics are also studied recently (Salamati2021DDCBC; wang2023stochastic; schon2024DRObarrier) with an available software tool LUCID (lucid). Despite these advances, most methods still depend on manual template selection and limited function classes, underscoring the need for adaptive, reasoning-driven synthesis frameworks.
LLMs for Complex Reasoning.
Recent advances in LLMs have demonstrated exceptional reasoning, planning, and problem-solving capabilities (achiam2023gpt; zhao2023survey). Models such as GPT-3 (dale2021gpt3) and GPT-4 (sanderson2023gpt4) exhibit strong generalization and inference abilities by leveraging billions of pretrained parameters and broad world knowledge (xu2024llm). LLMs can decompose complex problems into sub-tasks and reason across multiple abstraction levels, enabling coherent multi-step explanations and decisions (jansen2020visually; lin2023text2motion; rana2023sayplan). These capabilities have been applied across domains—from robotic task planning (rana2023sayplan) and autonomous driving (cui2023large) to mathematical reasoning tasks requiring symbolic manipulation, quantitative analysis, and formal logic (lewkowycz2022solving). Such progress positions LLMs as promising tools for safety-critical domains that demand rigorous mathematical reasoning about system constraints and safe execution.
LLM for CPS.
LLMs are increasingly transforming CPS by introducing high-level reasoning, planning, and natural-language interaction capabilities (xu2024llm). Beyond traditional machine learning models trained on narrow datasets, LLMs draw on diverse web-scale corpora of text and code to serve as both context-aware assistants and autonomous decision-making agents. In robotics, systems such as SayCan (ahn2022saycan) and RT-2 (zitkovich2023rt2) combine LLM reasoning with reinforcement learning or multimodal perception to translate natural-language commands into executable actions. In autonomous driving, DriveGPT4 (yang2023drivegpt4) and Talk2Drive (cui2023large) employ LLMs for interpretable decision-making and natural language interaction.
For formal controller synthesis, bayat2025llm propose a multi-agent framework that translates natural-language specifications into verified symbolic control code, combining reasoning with abstraction-based verification. Despite these advances, integrating LLMs into CPS remains challenging. LLMs are susceptible to hallucination (xu2024llm), lack grounding in physical constraints, and cannot be verified using traditional neural-network methods such as Reluplex due to scale and architectural complexity. These limitations motivate hybrid approaches that couple LLM reasoning with formal verification to ensure correctness and safety in real-world CPS.
6 Limitations
Although our synthesis framework outperforms the baselines by solving 90% of the BarrierBench problems, the remaining 10 unsolved problems reveal some synthesis challenges. The unsolved cases are highly complex nonlinear dynamics, including systems with complex combination of trigonometric and exponential terms, across different topological sets. Their complexity challenges both the synthesis framework and the verification solvers. In certain cases, the Synthesis Agent fails to generate barrier certificates satisfying all safety conditions due to the complexity of the underlying dynamics. Moreover, for the synthesized barriers, SMT solvers face computational difficulties during verifying conditions, frequently timing out or failing to determine satisfiability.
7 Conclusion
We present a novel LLM agentic framework for barrier certificate synthesis that addresses fundamental limitations in existing approaches through natural language reasoning and adaptive mathematical structure exploration. By leveraging the reasoning capabilities of LLMs augmented with relevant barrier retrieval and verification-guided refinement, our framework automates the traditionally manual process of template selection and enables the discovery of diverse barrier certificate structures that fixed-template approaches can not discover. Moreover, we extend the framework to simultaneous barrier-controller co-synthesis, ensuring mathematical compatibility between safety certificates and feedback control laws. We introduce BarrierBench, a benchmark of 100 dynamical systems spanning continuous-time, discrete-time, linear, nonlinear, and controlled systems across various dimensions. Our evaluation demonstrates 90% success rate of the framework on this benchmark.
This work presents an initial step toward integrating language-based reasoning with formal safety verification and control synthesis. While the proposed framework and BarrierBench benchmark provide a concrete foundation for exploring this direction, we do not claim completeness or exhaustive coverage of all system classes, synthesis strategies, or verification settings. Instead, our goal is to establish a starting point and an open, extensible database that the community can collectively expand and refine. By enabling shared experimentation and collaborative benchmarking, we hope to accelerate progress toward more general, interpretable, and automated methods for safety assurance in dynamical systems. The authors remain open to any suggestions, feedback, and contribution of new systems to further develop and enrich BarrierBench.
Appendix A LLM Prompts
This appendix presents the complete set of prompts used in our framework.
A.1 Prompt 1: Dataset Similarity Selection
This prompt selects the most similar problem from filtered candidates using LLM-based similarity assessment.
Dataset Similarity Selection TARGET PROBLEM:
Dynamics: {SYSTEM_DYNAMICS}
Initial set: {INITIAL_SET}
Unsafe set: {UNSAFE_SET}
COMPATIBLE CANDIDATES (all are fundamentally similar):
{CANDIDATES_TEXT}
Which candidate has the most similar problem type and structure to the target problem? Focus on: system structure, problem type, and mathematical pattern similarity. Answer with only the candidate number (1, 2, 3, etc.):
A.2 Prompt 2: Barrier Generation in the First Iteration
This prompt generates the barrier certificate for systems incorporating the retrieved context if available.
A.2.1 Systems Without Controller
Barrier Synthesis in the First Iteration - No Controller {CONTEXT} Main Problem:
- Dynamics: {SYSTEM_DYNAMICS}
- Initial set: {INITIAL_SET}
- Unsafe set: {UNSAFE_SET}
Design a barrier certificate B(x) that satisfies:
1. B(x) $\leq$ 0 in initial set
2. B(x) > 0 in unsafe set
3. {CONDITION_3}
Be very careful - don’t make it more complex than needed. CRITICAL: Use ONLY real numbers in the barrier expression. No variables like ’c’ or $\varepsilon$’. Solve specifically for THIS problem with appropriate coefficients. Analyze carefully but be concise. Give precise answer without long explanations. Format your response as (don’t make it bold): BARRIER: [expression with numbers only]
The variable CONDITION_3 refers to
  • Discrete-time: $B(f(x))-B(x)\leq 0$ for all $x$ in the state space;
  • Continuous-time: $\nabla B(x)\cdot f(x)<0$ on the boundary.
The variable CONTEXT refers to the following two cases:
(a) when a similar problem exists in the dataset:
Related problem found:
EXAMPLE: {SIMILAR_PROBLEM}
B(x): {SIMILAR_BARRIER}
WARNING: This is just an example - your solution may be
completely different NOT ONLY in terms of coefficients,
BUT ALSO in format and structure. You may make mistakes,
so do not consider this as a solution. Please dont copy it.
(b) when no similar problem exists:
No similar problems found. Analyze this problem fresh.
A.2.2 Systems With Control Inputs
Barrier and Controller Synthesis in the First Iteration {CONTEXT} Main Problem:
- Dynamics: {SYSTEM_DYNAMICS}
- Initial set: {INITIAL_SET}
- Unsafe set: {UNSAFE_SET}
CONTROLLER SYNTHESIS: This system has control inputs {CONTROLLER_PARAMETERS}. You need to design BOTH:
1. Barrier certificate B(x)
2. Controller expressions for {CONTROLLER_PARAMETERS}
The controller u(x) will be substituted into dynamics to create closed-loop system. Design a barrier certificate B(x) that satisfies:
1. B(x) $\leq$ 0 in initial set
2. B(x) > 0 in unsafe set
3. {CONDITION_3}
Be very careful - don’t make it more complex than needed. Design both barrier certificate B(x) AND controller expressions that work
together to satisfy all conditions.
CRITICAL:
- Use ONLY real numbers in both barrier and controller expressions.
No variables like ’c’ or $\varepsilon$’.
- Solve specifically for THIS problem with appropriate coefficients.
- Controller must be implementable with realistic actuators
- Ensure controller bounds are reasonable (avoid extremely large values)
Analyze carefully but be concise. Give precise answer without long explanations. Format your response as (don’t make it bold): BARRIER: [barrier expression with numbers only]
CONTROLLER: [controller expressions for each parameter, comma-separated]
A.3 Prompt 4: Barrier Generation in Subsequent Iterations
For subsequent iterations, the prompt leverages failure history to improve structure.
A.3.1 Systems Without Controller
Barrier Synthesis in Subsequent Iterations - No Controller Previous attempts failed:
- Tried: {BARRIER} {FAILED_INFO}
- Tried: {BARRIER} {FAILED_INFO}
Improve the barrier structure to satisfy all conditions. Main Problem:
- Dynamics: {SYSTEM_DYNAMICS}
- Initial set: {INITIAL_SET}
- Unsafe set: {UNSAFE_SET}
Design a barrier certificate B(x) that satisfies:
1. B(x) $\leq$ 0 in initial set
2. B(x) > 0 in unsafe set
3. {CONDITION_3}
Learn from previous failures. You can change structure of TEMPLATE if needed. In this step, the goal is to improve the structure of the templates, not refine the parameters.
CRITICAL: Use ONLY real numbers in the barrier expression. No variables like ’c’ or $\varepsilon$’. Solve specifically for THIS problem with appropriate coefficients. Analyze carefully but be concise. Give precise answer without long explanations. Format your response as (don’t make it bold): BARRIER: [expression with numbers only]
The variable FAILED_INFO refers to
failed: FAILED_CONDITIONS (number of counter-examples).
A.3.2 Systems With Control Inputs
Barrier and Controller Synthesis in Subsequent Iterations Previous barrier + controller attempts failed:
- Barrier: {BARRIER}, Controller: {CONTROLLER}, {FAILED_INFO}
- Barrier: {BARRIER}, Controller: {CONTROLLER}, {FAILED_INFO}
- Barrier: {BARRIER}, Controller: {CONTROLLER}, {FAILED_INFO}
Improve the barrier + controller structure to satisfy all conditions. Main Problem:
- Dynamics: {SYSTEM_DYNAMICS}
- Initial set: {INITIAL_SET}
- Unsafe set: {UNSAFE_SET}
CONTROLLER SYNTHESIS:
This system has control inputs {CONTROLLER_PARAMETERS}. You need to
design BOTH:
1. Barrier certificate B(x)
2. Controller expressions for {CONTROLLER_PARAMETERS}
The controller u(x) will be substituted into dynamics to create closed-loop system. Design a barrier certificate B(x) that satisfies:
1. B(x) $\leq$ 0 in initial set
2. B(x) > 0 in unsafe set
3. {CONDITION_3}
Learn from previous failures. You can change structure of TEMPLATE if needed. In this step, the goal is to improve the structure of the templates, not refine the parameters. Design both barrier certificate B(x) AND controller expressions that work
together to satisfy all conditions.
CRITICAL:
- Use ONLY real numbers in both barrier and controller expressions.
No variables like ’c’ or $\varepsilon$’.
- Solve specifically for THIS problem with appropriate coefficients.
- Controller must be implementable with realistic actuators
- Ensure controller bounds are reasonable (avoid extremely large values)
Analyze carefully but be concise. Give precise answer without long explanations. Format your response as (don’t make it bold): BARRIER: [barrier expression with numbers only]
CONTROLLER: [controller expressions for each parameter, comma-separated]
A.4 Prompt 6: Coefficient Refinement in the First Two Iterations
For the first two iterations of the refinement, the prompt focuses on the adjustment of the coefficient without structural changes.
A.4.1 Systems Without Controller
Coefficient Refinement - No Controller Original barrier: {BARRIER} {FAILED_INFO} {REFINEMENT_HISTORY} Problem:
- Dynamics: {SYSTEM_DYNAMICS}
- Initial set: {INITIAL_SET}
- Unsafe set: {UNSAFE_SET}
Try a different coefficient distribution. You can redistribute the coefficients between ALL terms (sometimes it is necessary for all terms to have different coefficients), but DO NOT change structure Requirements:
1. B(x) $\leq$ 0 in initial set
2. B(x) > 0 in unsafe set
3. {CONDITION_3}
Analyze carefully but be concise. Give precise answer without long explanations. Format your response as (don’t make it bold): REFINED_BARRIER: [expression with numbers only]
A.4.2 Systems With Control Inputs
Coefficient Refinement - Barrier and Controller Original barrier: {BARRIER}, Original controller: {CONTROLLER}, {FAILED_INFO} {REFINEMENT_HISTORY} Problem:
- Dynamics: {SYSTEM_DYNAMICS}
- Initial set: {INITIAL_SET}
- Unsafe set: {UNSAFE_SET}
Try a different coefficient distribution for both barrier and controller.
You can redistribute the coefficients between ALL terms, but DO NOT change structure
CONTROLLER SYNTHESIS CONSTRAINTS:
1. Controller parameters: {CONTROLLER_PARAMETERS}
2. Use smooth, bounded functions (avoid extremely large values)
3. Controller must work harmoniously with the barrier
4. Ensure closed-loop stability
Requirements:
1. B(x) $\leq$ 0 in initial set
2. B(x) > 0 in unsafe set
3. {CONDITION_3}
Analyze carefully but be concise. Give precise answer without long explanations. Format your response as (don’t make it bold): REFINED_BARRIER: [barrier expression with numbers only]
REFINED_CONTROLLER: [controller expressions for each parameter, comma-separated]
A.5 Prompt 7: Structure Refinement in Subsequent Iterations
For refinement iterations 3 and 4, the prompt allows structural modifications when coefficient adjustments fail.
A.5.1 Systems Without Controller
Structure Refinement - No Controller Original barrier: {BARRIER} {FAILED_INFO} {REFINEMENT_HISTORY} Problem:
- Dynamics: {SYSTEM_DYNAMICS}
- Initial set: {INITIAL_SET}
- Unsafe set: {UNSAFE_SET}
Previous coefficient adjustments failed. Consider changing the barrier structure if needed while keeping the same polynomial degree. You can modify the terms or their combinations. Requirements:
1. B(x) $\leq$ 0 in initial set
2. B(x) > 0 in unsafe set
3. {CONDITION_3}
Analyze carefully but be concise. Give precise answer without long explanations. Format your response as (don’t make it bold): REFINED_BARRIER: [expression with numbers only]
A.5.2 Systems With Control Input
Structure Refinement - Barrier and Controller Original barrier: {BARRIER}, Original controller: {CONTROLLER}, {FAILED_INFO} {REFINEMENT_HISTORY} Problem:
- Dynamics: {SYSTEM_DYNAMICS}
- Initial set: {INITIAL_SET}
- Unsafe set: {UNSAFE_SET}
Previous coefficient adjustments failed. Consider changing the barrier and/or controller structure if needed while keeping the same polynomial degree. You can modify the terms or their combinations. CONTROLLER SYNTHESIS CONSTRAINTS:
1. Controller parameters: {CONTROLLER_PARAMETERS}
2. Use smooth, bounded functions (avoid extremely large values)
3. Controller must work harmoniously with the barrier
4. Ensure closed-loop stability
Requirements:
1. B(x) $\leq$ 0 in initial set
2. B(x) > 0 in unsafe set
3. {CONDITION_3}
Analyze carefully but be concise. Give precise answer without long explanations. Format your response as (don’t make it bold): REFINED_BARRIER: [barrier expression with numbers only]
REFINED_CONTROLLER: [controller expressions for each parameter, comma-separated]
Refinement History Format:
Refinement 1: {BARRIER} {FAILED_INFO}
Refinement 2: {BARRIER} {FAILED_INFO}
A.6 Prompt 8: SMT Solver
A.6.1 Solver Selection
This prompt selects the most appropriate SMT solver for verification.
SMT Solver Selection Select the best SMT solver based on barrier expression and the dynamical system. PROBLEM:
Dynamics: {SYSTEM_DYNAMICS}
Barrier: {BARRIER_EXPRESSION}
AVAILABLE SOLVERS:
- cvc5
- z3
- yices
Without long explanations, give a short and precise answer. Format your response as: SOLVER: [solver name]
A.6.2 Timeout Analysis
This prompt determines whether to retry with extended timeout when a solver times out.
SMT Solver Timeout Analysis Solver {SOLVER_NAME} timed out after {TIMEOUT_MS}ms. PROBLEM:
Dynamics: {SYSTEM_DYNAMICS}
Barrier: {BARRIER_EXPRESSION}
Given the above information, should we attempt again with more time, or is this barrier too complex to verify? Without long explanations, give a short and precise answer. Format your response as: RETRY: yes or no
TIMEOUT_MULTIPLIER: [number] (only if RETRY is yes, e.g., 1.5 or 2.0)
A.6.3 Error Analysis
This prompt selects an alternative solver when verification fails.
Alternative Solver Selection Solver {SOLVER_NAME} failed during verification. ERROR: {ERROR_TYPE}
MESSAGE: {ERROR_MESSAGE}
PROBLEM:
Dynamics: {SYSTEM_DYNAMICS}
Barrier: {BARRIER_EXPRESSION}
REMAINING SOLVERS:
{REMAINING_SOLVERS_LIST}
Select a different solver to try. Without long explanations, give a short and precise answer. Format your response as: NEXT_SOLVER: [solver name]